OpenStudy (anonymous):

A candy store sells malted milk balls for $2.50 per pound, and chocolate-covered raisins for $3.75 per pound. How many pounds of each should be used to make a 10 pound mixture that sells for $3.00 per pound?

OpenStudy (anonymous):

m + r = 10 2.5m+3.75r=10(3) 2.5(10-r)+3.75r=30 1.25r+25=30 1.25r=5 r=4 m=6